the perimeter of a rectangle is 44 inches.Find the dimension if the width is two inches longer than three times the length.Then find the area of the rectangle
Answer by lwsshak3(11628)

the perimeter of a rectangle is 44 inches.Find the dimension if the width is two inches longer than three times the length.Then find the area of the rectangle.
**
let x=length of rectangle
3x+2=width of rectangle
Perimeter=2*(length+width)=44
=2(x+3x+2)=2(4x+2)=8x+4=44
8x=40
x=5 (length
3x+2=17 (width)
Area of rectangle=length*width=5*17=85 sq in